differentiate cos x with respect to e
 
 
another question is if ma=2i+3j+2k is a unit vector find value of m

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Please ask one question in one thread.
 
Now
1st Answer-
 
We have to find d(cos x) /d (e^x)
 
We can also find 
d cos x/ dx   and d (e^x)/ dx. And then divide.
 
Now
 
d ( cos x) /dx = - sinx 
 
d (e^x) / dx = e^x
 
Hence
 
d (cos x) / dx = - sinx /e^x
 
 
Regards
Arun (askIITians forum expert)